OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale

OpenStudy (anonymous):

1, x = 100 Plug in your equation, C. C = 0.765(100) + 0.06[0.765(100)] C = 76.5 + 0.06(76.5) C = 76.5 + 4.59 C = 81.09